# Proactive Agents for Multi-Turn Text-to-Image Generation Under Uncertainty

### v2

@@ 7.2 Automatic Metrics and Qualitative Results
Coco-Captions T2I 0.8757±.$\pm.03$ 0.5170±.$\pm.16$ 0.2976±.$\pm.45$ 520.0645±$\pm 161.3$ $\pm.05$ 0.1735
Ag1 0.9440±.$\pm.02$ 0.6269±.$\pm.12$ 0.5831±.$\pm.49$ 508.4014±$\pm 158.5$ 0.7555±.$\pm.08$ 0.2024
Ag2 0.9461±.$\pm.02$ 0.6141±.$\pm.13$ 0.6632±.$\pm.46$ 481.7224±$\pm 154.5$ $\pm.08$ 0.2326
Ag3 0.9501±.$\pm.02$ 0.6575±.$\pm.10$ 0.7751±.$\pm.39$ 446.5679±$\pm 151.8$ 0.9001±.$\pm.05$ 0.2915
ImageInWords T2I 0.8807±.$\pm.02$ 0.5154±.$\pm.15$ 0.3711±.$\pm.47$ 459.9053±$\pm 200.2$ 0.6815±.$\pm.70$
Ag1 0.9429±.$\pm.02$ 0.5548±.$\pm.15$ 0.5058±.$\pm.48$ 449.8927±$\pm 196.1$ $\pm.08$
Ag2 0.9382±.$\pm.02$ 0.5645±.$\pm.15$ 0.5701±.$\pm.48$ 444.5227±$\pm 193.7$ 0.8791±.$\pm.07$
Ag3 0.9418±.$\pm.02$ 0.5875±.$\pm.14$ 0.6624±.$\pm.45$ 429.4636±$\pm 194.5$ 0.9124±.$\pm.06$
DesignBench T2I 0.8740±.$\pm.02$ 0.5439±.$\pm.12$ 0.3528±.$\pm.48$ 320.8898±$\pm 93.7$ 0.6074±.$\pm.08$ 0.0818
Ag1 0.9365±.$\pm.02$ 0.5943±.$\pm.12$ 0.6848±.$\pm.46$ 295.1974±$\pm 69.2$ $\pm.08$ 0.1133
Ag2 0.9384±.$\pm.02$ 0.6417±.$\pm.11$ 0.8553±.$\pm.34$ 271.2604±$\pm 81.9$ 0.9181±.$\pm.06$ 0.2692
Ag3 0.9429±.$\pm.02$ 0.6924±.$\pm.12$ 0.9545±.$\pm.21$ 257.4352±$\pm 67.5$ 0.9485±.$\pm.04$ 0.5357
Table 1: Automatic evaluation results on Coco-Captions, ImageInWords, and DesignBench. Our agents show large performance gains in all metrics over a standard T2I model alone. Human Rating displays the % of times an image was ranked as closest to the ground truth image. All other columns correspond to automatic evaluations using metrics described in 7.1.3.
7.3 Human Studies on Generated Images and Dialogues
We conducted human studies to verify the automatic evaluations of the agents. See detailed design in Figure 19, Figure 20 and Figure 21. Information on human user recruitment and consent can be found in §H.
Table 1 shows the human rater evaluations on the generated images in comparison to the ground truth image. Participants are asked to rank the images produced by the three proposed multi-turn agents and a single-turn T2I model in terms of content and style against a ground truth image for which the original prompt was derived and the answers to the agent’s questions were derived. Approximately 550 image-dialog pairs per agent are rated using 3 human raters. The generated unlabeled images were presented in a random order and the human rater was tasked with ranking the images from best to worst in terms of content correctness. The results are in Table 1, under the column Human Rating, showing that agentic systems are selected as the best generated image over the single-turn T2I in 80%+ of cases for both content and style. Human ratings of the image directly correspond to and validate the I2I (DINO) results.
Figure 4 shows the human study results on the generated agent-simulated user dialogues. (1) Per question raters are asked to mark issues each agent question contains that could pose a disturbance to the user, i.e., ‘question is too long and complex’. Approximately 8k questions per agent are rated. The results in Figure 4 (left), show agentic systems have issues with their questions in 14% or less cases. Results show Ag2 and Ag3 suffer from too long questions while Ag1 suffers from questions that do not gain new information. (2) Raters are asked to evaluate the correspondence of each image to the agent-user dialog and original prompt. Approximately 1.5k image-dialog pairs are rated using 3 human raters. Results in Figure 4 (right) show that for all of our agents, more than 96% of the 1.5k image-dialog pairs are rated as very close or fairly close with some differences. This high rating shows the viability of the T2I model employed by the agents, as well as the agents’ ability to combine the dialogue into a coherent prompt to feed the T2I model.
Figure 4: Human study results on the dialogues between agents and simulated users. Left: The distribution of issues in agent questions - per agent, as determined by human raters. Right: The rating of how well the final generated image corresponds to the user prompt and dialogue.
@@ 7.4 Human Studies on the Agent Interface
Table 4 in Appendix confirms the prevalence of hypothesized frustrations amongst users, with 83% experiencing occasional, frequent, or very frequent frustration due to prompt iterations, followed by 70% for misinterpretations, 71% for inconsistent generations, and 60% experiencing frustration due to incorrect assumptions. Most acutely 55% of participants reported frequent or very frequent frustration due to the prompt iteration frequency necessary. In Table 2, we report the mitigation features that are likely to help. Clarifications were reported as having the highest likelihood to help current workflows (91% could / likely / very likely to be helpful), followed by entity graphs (88% could / likely / very likely to be helpful) and relationship graphs (86% could / likely / very likely to be helpful). Clarifications were expected to deliver value immediately / very soon by 58%.
@@ 8 Discussion and Conclusion
While T2I prompt-image alignment errors can limit the effectiveness of our proposed agents, our experiment in §E.2 demonstrates that by leveraging agent-user QA pairs to evaluate image fidelity, we can filter out misaligned images across N seeds, thereby mitigating the impact of these errors. Note that the agents’ T2T scores in Table 1 (92%+) ablate the T2I model, showing that they have already reached high alignment on the caption level, which is bound to increase with improved T2I capabilities.
@@ Acknowledgements
We would like to thank Jason Baldridge and Zoubin Ghahramani for insightful discussions on multi-turn T2I and belief states, Mahima Pushkarna for the help and consultation on user study. We would also like to thank Richard Song, Noah Fiedel and anonymous reviewers for feedback on the paper.
@@ Appendix B Formalism of the Agent and its Objective
  • $A$: a space of actions that the agent can take,
@@ Appendix C Visualization of Multi-Turn Agent-User Dialogs and Generated Images
In Figure 5, we show examples of multi-turn dialogs between simulated users and the three agents in Section 7. We also visualize the generated images in Figure 6, Figure 7, Figure 8 and Figure 9.
@@ D.1 Implementation of Agent Beliefs
Belief Parsing. See an example of the belief graph in fig. 10. We employ three specialized parsers trained via in-context learning (ICL): entity parser (§D.5) analyzes the user prompt to identify and extract a list of relevant entities.; attribute parser (§D.6) takes user prompt and an entity as the input to extract a list of attributes associated with that entity; relation parser (§D.7) takes the user prompt and a list of entities as input and identifies relationships between those entities. Each entity is associated with meta information like name, importance to ask score, description, probability of appearing, a list of attributes like color, position, etc  [^6] Name is a unique identifier for the entity; Importance to ask score: A numerical value indicating the entity’s perceived importance in satisfying the user’s request. Entities with higher scores are prioritized during question generation, as they are likely to reduce uncertainty and contribute significantly to the final image; Description provides a textual description of the entity; probability of appearing estimates likelihood of the entity being present in the generated image; Attributes are for understanding the detailed attributes of the entities.. Each attribute contains meta information like name, importance to ask score, a list of possible values for the attribute along with their associated probabilities, etc. Each relation includes meta information such as: name, description, spatial relation, importance to ask score, entity 1 and entity 2, whether the relation is bidirectional, etc.
Figure 10: An example of the belief graph data structure for a given prompt in Figure 1.
@@ Appendix E Automatic Evaluation
In Algorithm 2, we show the user-agent self-play procedures that we used to perform all automated evaluation.
@@ E.1 DSG Evaluations
The DSG (Cho et al., 2024) metric is used to compare the ground truth caption and the generated caption. DSG is adapted to parse text prompts into Davidsonian scene graph using the released code. The plots in Figure 11 show the T2T DSG metrics per turn. The results in Figure 11 show that significant gains in performance come from using proactive multi-turn agents. The blue line on the graph represents the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt p0. Note that all multi-turn agents far exceed the baseline T2I model on all benchmarks for the DSG metric.
@@ E.2 Mitigation of T2I Errors
1:  Input: Initial prompt $p_{0}$, User $u$, Agent $a$ (with $p_{0}$), $max\_turns$
@@
3:  $p_{f}\leftarrow p_{0}$
4:  for $turn\_id=0$ to $max\_turns-1$ do
@@ 1. Default state
On load of the app, there would be a text prompt input and space for output images, as is common across typical T2I interfaces. There would also be space for the user to view either clarifications from the model, or a graph interface, as part of the overall “input” section as these would act as a further input for future model output iterations. See Figure 12 below as reference.
@@ 2. Output images, with Clarifications
Once the user has submitted the prompt and the model has responded, there would be a set of images, as initial outputs from the users prompt. Below the input prompt would be a set of “Clarifications” in its populated state. These clarifications would ask the user specific questions that would be necessary to increase the specificity of the prompt, for the model to get a more accurate results aligned to the users intention, or to help the user realise their intention. Options would be given of the highest probability options for each Clarification, but the user could also fill in a totally new option via a free text field. Once answered by selection or text input, the clarifications would be added to the above, primary prompt for regeneration when the user selects. See Figure 13 below as reference.
@@ 3. Graph Entities & Attributes
Below the Graph would also be a list of ’cards’ (i.e. boxed groups of information), one for each "explicit" or "implicit" entity. Within each card a user could see the status of implicit / explicit, and change this status to confirm or deny its presence. The user could also see a list of "attributes" associated to that entity, which the model has assumed. Each of these attributes could be changed by interacting with a list of alternatives via drop down. These lists are determined in terms of which items and order of items, based on the probability by which the model sees them, ordered with highest first. This probability would be made clear to the user to define the order by seeing the percentage next to the label. See Figure 14 below as reference.
@@ 4. Graph Relationships
The user would also be able to change the state of the Graph and Cards, to instead focus on the relationships between entities, by toggling to "Relations". In this state the user would be able to focus on two specific entities (e.g. ’mug’ and ’table’), see the description of the relationship (e.g. ’the mug is sitting on the table’) and if desired change the relationship to an alternative (e.g. ’on’, changed to ’under’) via a drop down of options which the model determined as alternative options ordered by probability, as per attributes. See Figure Figure 15 below as reference.
@@ G.3 Human Study Results
Table 3 details the T2I usage frequency of the human subjects. Table 4 shows the percentage of human subjects that reported different kinds of frustrations in their experience of using T2I. Table 5 summarizes the results on expected speed of value delivered from different features of our agent prototypes. These results highlight the impact of our contributions.
@@
Table 4: Reported User Frustrations with existing T2I processes (of participants)
@@
Table 5: Expected speed of value delivered from features (of users)
@@ Appendix H Details on Recruitment and Participant Consent
Figure 19: An example of the template presented to human raters. Human raters are asked to mark any issues a question contains that could pose a disturbance to the user. Approximately 8k questions per Agent are rated. The results are shown in Figure 4.
@@
Figure 20: An example of the template presented to human raters. Human raters are asked to rank the correspondence of each image to the agent-user dialog and original prompt. Approximately 1.5k image-dialog pairs are rated using 3 human raters. Results in Figure 4.
@@
Figure 21: A template of the task presented to human raters. Human raters are asked to rate the images produced by the three proposed multi-turn agents and a single-turn T2I model against a ground truth image for which the original prompt was derived and the answers to the agents questions were derived. Approximately 550 image-dialog pairs per agent are rated using 3 human raters. The generated images were presented in a random order and were unlabeled and the human rater was tasked with ranking the images from best to worst. The results from the study are shown in Figure 4.

### v3

@@ 7.3 Human Studies on Generated Images and Dialogues
[…]
Table 1: Automatic evaluation results on Coco-Captions, ImageInWords, and DesignBench. Our agents show large performance gains in all metrics over a standard T2I model. The Human Eval columns display the % of times an image was ranked as closest to the ground truth image in terms of content or aesthetics. All other columns correspond to automatic evaluations using metrics described in 7.1.3.
Figure 4: Human study results on the dialogues between agents and simulated users. (a) Issues in each agent’s questions, as determined by human raters. (b) Ratings of how well the final generated image corresponds to the user prompt and dialogue.
@@ 7.4 Human Studies on the Agent Interface
Table˜4 in Appendix confirms the prevalence of hypothesized frustrations amongst users, with 83% experiencing occasional, frequent, or very frequent frustration due to prompt iterations, followed by 70% for misinterpretations, 71% for inconsistent generations, and 60% experiencing frustration due to incorrect assumptions. Most acutely 55% of participants reported frequent or very frequent frustration due to the prompt iteration frequency necessary. In Table˜2, we report the mitigation features that are likely to help. Clarifications were reported as having the highest likelihood to help current workflows (91% could / likely / very likely to be helpful), followed by entity graphs (88% could / likely / very likely to be helpful) and relationship graphs (86% could / likely / very likely to be helpful). Clarifications were expected to deliver value immediately / very soon by 58%.
@@ 8 Discussion and Conclusion
While T2I prompt-image alignment errors can limit the effectiveness of our proposed agents, our experiment in §E.2 demonstrates that by leveraging agent-user QA pairs to evaluate image fidelity, we can filter out misaligned images across N seeds, thereby mitigating the impact of these errors. Note that the agents’ T2T scores in Table˜1 (92%+) ablate the T2I model, showing that they have already reached high alignment on the caption level, which is bound to increase with improved T2I capabilities.
@@ Acknowledgements
We would like to thank Jason Baldridge and Zoubin Ghahramani for insightful discussions on multi-turn T2I and belief states, Mahima Pushkarna for the help and consultation on user study. We would also like to thank Richard Song, Noah Fiedel, Susanna Ricco and anonymous reviewers for feedback on the paper.
@@ Appendix B Formalism of the Agent and its Objective
  • AA: a space of actions that the agent can take,
@@ Appendix C Visualization of Multi-Turn Agent-User Dialogs and Generated Images
In Figure˜5, we show examples of multi-turn dialogs between simulated users and the three agents in Section˜7. We also visualize the generated images in Figure˜6, Figure˜7, Figure˜8 and Figure˜9.
@@ D.1 Implementation of Agent Beliefs
Belief Parsing. See an example of the belief graph in fig.˜10. We employ three specialized parsers trained via in-context learning (ICL): entity parser (§D.5) analyzes the user prompt to identify and extract a list of relevant entities.; attribute parser (§D.6) takes user prompt and an entity as the input to extract a list of attributes associated with that entity; relation parser (§D.7) takes the user prompt and a list of entities as input and identifies relationships between those entities. Each entity is associated with meta information like name, importance to ask score, description, probability of appearing, a list of attributes like color, position, etc  [^7] Name is a unique identifier for the entity; Importance to ask score: A numerical value indicating the entity’s perceived importance in satisfying the user’s request. Entities with higher scores are prioritized during question generation, as they are likely to reduce uncertainty and contribute significantly to the final image; Description provides a textual description of the entity; probability of appearing estimates likelihood of the entity being present in the generated image; Attributes are for understanding the detailed attributes of the entities.. Each attribute contains meta information like name, importance to ask score, a list of possible values for the attribute along with their associated probabilities, etc. Each relation includes meta information such as: name, description, spatial relation, importance to ask score, entity 1 and entity 2, whether the relation is bidirectional, etc.
Figure 10: An example of the belief graph data structure for a given prompt in Figure˜1.
@@ Appendix E Automatic Evaluation
In Algorithm˜2, we show the user-agent self-play procedures that we used to perform all automated evaluation.
@@ E.1 DSG Evaluations
The DSG (Cho et al., 2024) metric is used to compare the ground truth caption and the generated caption. DSG is adapted to parse text prompts into Davidsonian scene graph using the released code. The plots in Figure˜11 show the T2T DSG metrics per turn. The results in Figure˜11 show that significant gains in performance come from using proactive multi-turn agents. The blue line on the graph represents the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt p0. Note that all multi-turn agents far exceed the baseline T2I model on all benchmarks for the DSG metric.
@@ E.2 Mitigation of T2I Errors
1:Input: Initial prompt $p_{0}$, User $u$, Agent $a$ (with $p_{0}$), max_turnsmax\_turns
@@
[…]
4:for turn_id=0turn\_id=0 to max_turns1max\_turns-1 do
@@ 1. Default state
On load of the app, there would be a text prompt input and space for output images, as is common across typical T2I interfaces. There would also be space for the user to view either clarifications from the model, or a graph interface, as part of the overall “input” section as these would act as a further input for future model output iterations. See Figure˜12 below as reference.
@@ 2. Output images, with Clarifications
Once the user has submitted the prompt and the model has responded, there would be a set of images, as initial outputs from the users prompt. Below the input prompt would be a set of “Clarifications” in its populated state. These clarifications would ask the user specific questions that would be necessary to increase the specificity of the prompt, for the model to get a more accurate results aligned to the users intention, or to help the user realise their intention. Options would be given of the highest probability options for each Clarification, but the user could also fill in a totally new option via a free text field. Once answered by selection or text input, the clarifications would be added to the above, primary prompt for regeneration when the user selects. See Figure˜13 below as reference.
@@ 3. Graph Entities & Attributes
Below the Graph would also be a list of ’cards’ (i.e. boxed groups of information), one for each "explicit" or "implicit" entity. Within each card a user could see the status of implicit / explicit, and change this status to confirm or deny its presence. The user could also see a list of "attributes" associated to that entity, which the model has assumed. Each of these attributes could be changed by interacting with a list of alternatives via drop down. These lists are determined in terms of which items and order of items, based on the probability by which the model sees them, ordered with highest first. This probability would be made clear to the user to define the order by seeing the percentage next to the label. See Figure˜14 below as reference.
@@ 4. Graph Relationships
The user would also be able to change the state of the Graph and Cards, to instead focus on the relationships between entities, by toggling to "Relations". In this state the user would be able to focus on two specific entities (e.g. ’mug’ and ’table’), see the description of the relationship (e.g. ’the mug is sitting on the table’) and if desired change the relationship to an alternative (e.g. ’on’, changed to ’under’) via a drop down of options which the model determined as alternative options ordered by probability, as per attributes. See Figure Figure˜15 below as reference.
@@ G.3 Human Study Results
Table˜3 details the T2I usage frequency of the human subjects. Table˜4 shows the percentage of human subjects that reported different kinds of frustrations in their experience of using T2I. Table˜5 summarizes the results on expected speed of value delivered from different features of our agent prototypes. These results highlight the impact of our contributions.
@@
Table 4: Reported User Frustrations with existing T2I processes (% of participants)
@@
Table 5: Expected speed of value delivered from features (% of users)
@@ Appendix H Details on Recruitment and Participant Consent
Figure 19: An example of the template presented to human raters. Human raters are asked to mark any issues a question contains that could pose a disturbance to the user. Approximately 8k questions per Agent are rated. The results are shown in Figure˜4.
@@
Figure 20: An example of the template presented to human raters. Human raters are asked to rank the correspondence of each image to the agent-user dialog and original prompt. Approximately 1.5k image-dialog pairs are rated using 3 human raters. Results in Figure˜4.
@@
Figure 21: A template of the task presented to human raters. Human raters are asked to rate the images produced by the three proposed multi-turn agents and a single-turn T2I model against a ground truth image for which the original prompt was derived and the answers to the agents questions were derived. Approximately 550 image-dialog pairs (250 from ImageInWords, 250 from COCO-Captions and 21 from DesignBench) per agent are rated using 3 human raters. The generated images were presented in a random order and were unlabeled and the human rater was tasked with ranking the images from best to worst. The results from the study are shown in Table˜1.